SLAC PEP-II
BABAR
SLAC<->RAL
Babar logo
HEPIC E,S & H Databases PDG HEP preprints
Organization Detector Computing Physics Documentation
Personnel Glossary Sitemap Search Hypernews
Unwrap page!
Det. Search
Who's who?
Meetings
FAQ
Images
Archive
Systems
Performance
Intern. region
Vertex Tracker
Drift chamber
DIRC
Calorimeter
IFR
LST
Magnet
Electronics
Trigger
Operations
Run Coordination
Contact Experts
Shift Takers Info
Operations Manual
Electronic Logbook
Ops Hypernews
Shift Signup
Check this page for HTML 4.01 Transitional compliance with the
W3C Validator
(More checks...)

Minutes of the Trigger Operation Meeting (June 13, 2006)

Attendees: Rainer Bartoldus, Al Eisner, Fang Fang, Arik Kreisel, Selina Li, Rahmat Rahmat, Dong Su, Dan Walker

Operations

  • We have been taking high lumi data although there was some interuptions due to a leak at the arc5 chamber last week.
  • TRG has no operational problems for the last week.
  • Next week's TRG operation meeting will be replaced by the TRG parallel session in the BaBar collaboration meeting.

EMT

  • The deadtime due to EMT has dropped substantially since Christ O'Grady fixed the ROM0 bug. We have a similar discussion as the May 30th meeting. There was a suggestion to get the EMT prescale set to 2 if not back to 1 since shipping out the EMT data doesn't seem to be the problem for deadtime anymore.

DCT/DCZ

  • Su Dong and Selina started to disassemble the old DCT crates on June 13th (Tuesday morning). See Su Dong's HN posting. The online code/setup will need modifications so that we can still test the TSF firmware for the new DCZ system against the old TSF's without BLT/PTD. During the next downtime, we might want to test taking comics with the TSF systems without BLT/PTD for the old DCT to see if this is possible. Su Dong pointed out that it might be possible to test the Y firmware using the X crate. The TSF-Y fiber splitter will sit on top of the X crate till the summer. After the firmware is completely done, we would want some fiber barriers to connect to the new DCZ unless the fibers are long enough to reach the DCZ crates, rather than still using the fiber splitter like now. This will be done during the summer shutdown.
  • We also decided to move the DCZ ROMs to the current location of the DCT ROMs during the summer shutdown. The crate mask 0x40000000 for DCZ should probably also changed to 0x800 to replace the old DCT mask. This is one of a few items in run control that we will modify during the summer down. A complete list of the summer TRG activities will be outlined for next week's collaboration meeting.

Online

  • Another online release is being built and it will include the DCZ online lib with teststand dependency removed.
  • Rainer pointed out that we need a L3 validation with the new online release since L3 had quite a lot of substantial changes went in. We don't know yet when the new online release will be installed at IR2 for data-taking.
  • Fang is working on the DCT calibration validation.
  • Another change proposed: filter the ZPD damage so that events with ZPD damage won't be in the all-event collection. This is a reminder for Rainer on his to-do list.

L3

  • Arik is working on track table for the L3 code.

Data Quality and Monitoring

  • Rahmat showed us a plot of the lumi ratio for bhabha to muon pair. The runs are grouped together for about 250/pb for each bin. The error bars still need verification. Al suggested that we should get this done and moved to production as soon as possible.

TrgConfig

  • BaBar's management would like to know what the prospect for the L1 trigger rate reduction is. The plan is to come out a tool for people who do analysis to run on their signal MC samples to see what sets of trigger lines fire. The results would help us find ways to tighten up the trigger configuration. Rainer will set up a module for this purpose to fill a 2-D histogram for trigger names correlational plots. We would like to have this tool ready and pilot tested on a few physics analyses by the collaboartion meeting. Before the next run period starts, we should be done with the tool and well tested with physics input from the collaboration.
  • Before the end of Run 5b, we will test the trigger rate by adding the 1Zn ("n" for negative tracks) which only involves the ZPD configuration. We will use the existing GLT object Zk and replace it with Zn.

This page is created by Selina Li (selina@slac.stanford.edu); however, you are welcome to edit and/or correct above items, specially for those who attended the meeting. Thanks.